OpenStudy (karatechopper):

is 12x=7y-10y a linear equation??

OpenStudy (turingtest):

12x=7y-10y 12x=(7-10)y=-3y y=-4x since this is of order 1 (i.e. it has no x^2, x^3, etc) it is linear